### 04 Jan 2024 v2

# perceptual musical features for interpretable audio tagging

Abstract
In the age of music streaming platforms, the task of automatically tagging music audio has garnered significant attention, driving researchers to devise methods aimed at enhancing performance metrics on standard datasets. Most recent approaches rely on deep neural networks, which, despite their impressive performance, possess opacity, making it challenging to elucidate their output for a given input. While the issue of interpretability has been emphasized in other fields like medicine, it has not received attention in music-related tasks. In this study, we explored the relevance of interpretability in the context of automatic music tagging. We constructed a workflow that incorporates three different information extraction techniques: a) leveraging symbolic knowledge, b) utilizing auxiliary deep neural networks, and c) employing signal processing to extract perceptual features from audio files. These features were subsequently used to train an interpretable machine-learning model for tag prediction. We conducted experiments on two datasets, namely the MTG-Jamendo dataset and the GTZAN dataset. Our method surpassed the performance of baseline models in both tasks and, in certain instances, demonstrated competitiveness with the current state-of-the-art. We conclude that there are use cases where the deterioration in performance is outweighed by the value of interpretability.
Index Terms—  Perceptual features, Explainable artificial intelligence, Music audio tagging
1 Introduction
Music audio tagging encompasses a range of tasks centered on assigning descriptive tags to music tracks, offering particular utility in the context of organizing extensive music collections. These tags encompass musically relevant information such as instruments [1], genres [2], mood [3], emotions [4, 5], harmony [6], rhythm [7], and additional metadata [8]. The majority of methods employed to address these tasks involve feature extraction and machine learning (ML). However, there has been a shift towards minimizing the feature extraction step, often limited to acquiring only Mel-frequency cepstrum coefficients (MFCCs), with a growing inclination towards end-to-end deep learning models [9]. This trend towards deep learning models brings with it increased opacity and challenges in providing explanations, particularly in scenarios where the ground truth itself can be subjective and open to interpretation [10, 11], as is the case with mood tags.
In these cases, the reliability of the training data dictates the reliability of the model, and without interpretability, it is difficult to detect flaws or biases in either data or models [12]. For music especially, where real-world data tend to be diverse, contrary to available datasets [13], it is essential to be able to detect and fix biases to make robust systems that can be applied in real-world settings [14]. As an extreme example, consider a mood tagging system that was trained only on pop and rock music. We cannot expect such a system to be generalized for all music genres, and it would not be possible to predict how it would behave if it were fed for instance classical music. Furthermore, we cannot expect an end-user to know the distribution of the training dataset, and the nuances of training machine learning models, thus such a system would not be applicable in practice, regardless of how well it performed on a test set. Contrarily, if the mood tagging system itself were interpretable, it would be easier for a user to understand why the system assigns specific tags to samples, making it more trustworthy and more likely to be applied in a real-world setting.
Fig. 1: Overview of the proposed pipeline
In recent years, efforts have emerged aimed at addressing the interpretability challenge within music audio tagging systems [15]. Perceptual musical features are used that are readily comprehensible to human users [16, 17, 18, 19, 20]. These features encompass characteristics imbued with both musical and acoustic significance, such as tonality, rhythmic stability, or loudness, and they exhibit a strong correlation with emotions [21, 22]. Due to this inherent correlation, these features have demonstrated their efficacy in Music Emotion Recognition (MER) tasks. These characteristics are rooted in musical concepts and theoretical knowledge, as noted in a survey [23], though quantitative studies on them remain relatively scarce. In our research, we build upon these concepts and execute a structured approach for achieving interpretable music audio tagging (depicted in Fig. 1), harnessing the power of perceptual features. Our specific contributions can be succinctly summarized as follows:
  • We incorporate three methods for extracting perceptual musical features including signal processing, deep neural networks, and symbolic knowledge.
  • We introduce a novel set of features related to music harmony by setting up a music knowledge base and using reasoning.
  • We interpret model predictions, using four different methodologies, namely SHAP [24], XGBoost [25] feature importance, permutation-based feature importance, and an ablation study.
  • Utilizing the interpretability methods, we discuss how we could alleviate issues with ML-based audio tagging, such as detecting biases of the training dataset which would be obscured in a deep learning setup, and better comprehend ambiguous labels.
2 Extracting perceptual features
2.1 Datasets
The dataset employed in our experiments was the MTG-Jamendo dataset [26], which serves as an open resource for music audio tagging. This dataset was constructed using music accessible through the Jamendo platform, offered under Creative Commons licenses, and enriched with tags provided by content uploaders. Boasting a substantial scale, it encompasses more than 18,486 complete audio tracks, annotated with 56 tags that encompass various mood and theme categories. We extracted MFCCs with 40 coefficients, utilizing a window and FFT size of 2048. Our choice of this specific dataset was predicated on the anticipation that interpretability would offer heightened value in a task involving partially subjective aspects, such as mood detection [27].
We utilized the GTZAN dataset [28] as our secondary dataset. This collection comprises 1000 audio tracks of 30 seconds and features 10 genres, with 100 tracks representing each. The audio files are all 16-bit monophonic recordings in .au format, sampled at 22kHz. We opted to use this dataset to assess the generalizability of our proposed pipeline on a smaller scale and a different corpus.
2.2 Methodologies
To extract perceptual features from the datasets, we employed three different techniques: Symbolic Knowledge, Deep Neural Networks, and Signal Processing.
By utilizing symbolic knowledge, we created a new set of interpretable features that are derived from music theory. To accomplish this, we constructed a knowledge base in order to extract music features from each track’s chords. We selected Omnizart [29] from among many chord recognition models, based on its reported performance. Our knowledge base consists of chords as instances, and classes and subclasses from the Functional Harmony Ontology (FHO) [30]. These classes and subclasses describe the harmonic function of each chord in a music track. Using SPARQL queries, we retrieved the instances of certain classes to create useful features, such as the frequency of specific harmonic patterns (e.g., dominant-dominant-tonic), in the form of n-grams.
We employed deep neural networks (DNNs) to efficiently capture perceptual features. A VGG-style model was trained on a dataset titled ’Mid-level perceptual musical features’  [17]. For input, we used Mel-Frequency Cepstral Coefficients (MFCCs) extracted from the first 30 seconds of each audio file yielded a matrix $K\in\mathbb{R}^{S\times P}$, where $S$ is the number of segments and $P$ is the count of MFCCs. Subsequently, the trained model was applied to infer perceptual features from our target datasets (MTG-Jamendo, GTZAN), facilitating a comprehensive analysis and extraction of perceptual features from the music audio data.
We employed signal processing techniques to extract acoustically and musically meaningful features. These features were thoughtfully curated by music experts with the specific criterion of ensuring their accessibility and comprehensibility to human perception. If $x$ is the input audio file, we define our process as a function $E$, where $E(x)=x_{s}$, $x_{s}\in\mathbb{R}^{q}$ and $q\in\mathbb{N}$ ($q$ being the number of the extracted features).
Fig. 2: Feature importance with two different methods
3 Experiments
When dealing with a set of perceptual features, there exist two strategies for achieving explainable predictions of audio tags. The first approach involves employing explainable-by-design machine learning models , like Decision Trees. The second approach revolves around the utilization of post hoc black-box explanation techniques capable of elucidating the functioning of any machine learning model [31]. In our research, we adopt a combination of both approaches. We incorporate explainable-by-design models due to their capacity to provide readily understandable outputs when fed with human-interpretable features as input. Additionally, we leverage post hoc black-box explanation methods, such as SHAP values, to enhance our interpretability efforts. We provide the GitHub repository for the implementation [^1] https://github.com/vaslyb/PerceptibleMusicTagging.
3.1 Classification
In our quest to select a classifier, we conducted a thorough evaluation of various machine-learning models using the MTG-Jamendo dataset. We conducted experiments with six classifiers, namely the Decision tree, K-Nearest Neighbors algorithm, Extra Trees Classifier, Naive Bayes classifier, Random Forest, and XGBoost. These classifiers were configured with their default hyper-parameter settings, and they were fed with the 62 scalar perceptual features as input. Our evaluation led us to choose XGBoost as the most effective classifier. XGBoost is an ensemble machine learning classifier based on decision trees that employs a gradient-boosting framework.
Our XGBoost model achieved a ROC-AUC score of 0.729, slightly surpassing the baseline of 0.725 while falling slightly below the performance of state-of-the-art intelligent systems (0.83) [32]. Similar results were obtained when applying XGBoost to the multi-class classification task using the GTZAN dataset, where we achieved an accuracy of 0.79, showing a slight deviation from the state-of-the-art performance (0.84) [33, 34].
While our findings may not outperform the state-of-the-art deep learning models, the presence of interpretable model outputs serves the valuable purpose of identifying potential biases within the model. Furthermore, we achieve commendable results using models characterized by a limited number of trainable parameters.
3.2 Explanations
In this study, we chose XGBoost as the predictive modeling technique due to its interpretability. It enables us to gain insights into the significance of input features in the overall predictions. This characteristic has rendered it a valuable tool for improving our understanding of the classification process.
Overall, we utilized four different methods to extract feature importance. Firstly, we used the model’s built-in feature importance. Secondly, we computed a set of feature importance explanations using the permutation method, by randomly shuffling each feature and computing the change in performance. Thirdly, we computed SHAP values, which estimate how each feature contributes to the prediction based on notions from game theory. Lastly, we conducted ablation studies, as recommended in the literature [35], with different groups of features. We categorized our features into six groups based on their origin - midlevel, signal, and harmonic - and conducted experiments with all possible combinations of the groups, using XGBoost as our classifier.
Fig. 3: Ablation studies. The above figure for the MTG Jamendo dataset and the below figure for the GTZAN dataset
Fig. 4: Interpreting label Disco of the GTZAN dataset
4 Discussion
The results of permutation-based and the overall weight XGBoost feature importance for the Jamendo dataset are shown in Fig. 2. The two methods produce similar results, which indicates that the displayed features are likely indeed important. Signal processing features were the most important for predictions, notably Spectral-Flux, Spectral-Complexity and Spectral-Centroid.
We obtained similar results in our ablation study (Fig. 3), where we found that signal processing features were the most useful, while harmonic features were the least useful for both datasets. Considering that harmonic features are extracted from the chord progression, this observation could be attributed to the quality of the chord recognition tool. Additionally, for the GTZAN dataset, only 30 sec for each track were given, leading to a limited number of chords per track and resulting in not having enough structural information. Harmonic features had a substantial impact only on predicting the Blues labels, a genre where chord progressions are distinct and easy to detect throughout a track. In terms of mid-level features, they appeared to be significant in improving the ROC-AUC metric for the Jamendo dataset. For both datasets, the best results were achieved by the combination of all features.
Fig. 5: Interpreting label Trailer of the MTG Jamendo dataset
We show examples of interpreting feature importance computed with SHAP values and XGBoost for predicting mood labels in the MTG Jamendo dataset. Specifically, we analyze one of the best-performing labels (ROC-AUC > 0.8) as shown in Fig. 5. For the highly-predictable Trailer mood predictions rely on low values of Length and high values of Spectral-Flux. A feature that is used for voice detection. This is aligned with the fact that trailers are usually shorter than songs and have voice-overs.
For GTZAN, we show examples of the interpretation for one predicted genre, Disco that reached 0.72 F1-score (see Fig. 4). For this label, it is noticeable that high values of Beat-Loud and middle values of BPM are the most prevalent features. These observations align with the common perception of Disco music as a genre characterized by distinguishable beats, easy to dance to.
5 Conclusions and future work
We have created a specialized pipeline for music audio tagging with an emphasis on delivering results that are interpretable. Our approach incorporates established feature extraction methods alongside the introduction of a fresh set of features derived from the realm of music harmony concepts. Our experimental findings underscore the pivotal role of feature engineering within our proposed pipeline, as it exerts an influence on both performance and interpretability. It is imperative that each feature’s meaning remains comprehensible to humans, emphasizing the significance of this step.
For future work, we plan to further investigate ways for improving performance given the explanations of predictions. Furthermore, we plan to study explanations regarding not only their performance improvement capabilities but also their usefulness and informativeness for end-users of music audio tagging systems. Finally, we plan to extend the feature extraction procedure with more features that are based on notions of music theory.